The top of the pool table is 0.870 m from the floor. The placement of the tape is such that 0 m is aligned with the edge of the table (as shown). The winner of the competition wants to know if he has broken the world record for the break shot of 32 mph (about 14.3 m/s). If the winners ball landed a distance 4.85 m from the table edge, calculate his break shot speed. Number m/s At what speed did his pool ball hit the ground Number m/s
Answer #1

along vertical

initial velocity voy = 0


acceleration ay = -g = -9.8 m/s^2

displacment y = -0.87 m


from equation of motion


y = voy*t + (1/2)*ay*t^2


-0.87 = 0 -(1/2)*9.8*t^2


t = 0.421 s


along horizontal


break speed vx = d/t = 4.85/0.421 = 11.5 m/s


=======================

after hitting the ground


vertical componnet vy = voy + ay*t


vy = 0 - 9.8*0.421 = -4.13 m/s


speed = v= sqrt(vy^2+vx^2) = sqrt(11.5^2+4.13^2) = 12.2 m/s